# Class-Independent Increment: An Efficient Approach for Multi-label Class-Incremental Learning

### v1

@@ Abstract
Current research on class-incremental learning primarily focuses on single-label classification tasks. However, real-world applications often involve multi-label scenarios, such as image retrieval and medical imaging. Therefore, this paper focuses on the challenging yet practical multi-label class-incremental learning (MLCIL) problem. In addition to the challenge of catastrophic forgetting, MLCIL encounters issues related to feature confusion, encompassing inter-session and intra-feature confusion. To address these problems, we propose a novel MLCIL approach called class-independent increment (CLIN). Specifically, in contrast to existing methods that extract image-level features, we propose a class-independent incremental network (CINet) to extract multiple class-level embeddings for multi-label samples. It learns and preserves the knowledge of different classes by constructing class-specific tokens. On this basis, we develop two novel loss functions, optimizing the learning of class-specific tokens and class-level embeddings, respectively. These losses aim to distinguish between new and old classes, further alleviating the problem of feature confusion. Extensive experiments on MS-COCO and PASCAL VOC datasets demonstrate the effectiveness of our method for improving recognition performance and mitigating forgetting on various MLCIL tasks.
Index Terms:
Multi-label class-incremental learning, Feature confusion, CINet, Class-level embedding, Contrastive loss.
I Introduction
@@
In the past decade, deep learning has catalyzed significant breakthroughs in computer vision and multimedia tasks, including image classification [1, 2], semantic segmentation [3], and object detection [4, 5], among others. However, the deep learning-based methods fail to recognize new samples of unseen classes. Directly learning these new samples will lead to catastrophic forgetting (CF), where the model’s performance on historical data significantly deteriorates. In this case, class-incremental learning (CIL) [6, 7, 8, 9, 10] is extensively studied to update the model only with new class data in each task while preserving the knowledge of old classes. However, in practical applications, an image typically contains multiple objects. For example, photos on social media often contain multiple labels or objects, making traditional CIL methods that assume each image contains only one object unsuitable. Therefore, some work explores studying CIL in the multi-label scenario, named multi-label class-incremental learning (MLCIL) [11, 12, 13, 14]. It aims to gradually learn new classes from label-incomplete multi-label data streams in different learning stages while maintaining performance on old labels.
@@
Compared to the traditional (i.e. single-label) CIL problem, it is more challenging as a single image may contain multiple objects of both the old and new classes, with only new class objects (one or multiple categories) annotated in each learning session. To this end, prior methods to address forgetting for MLCIL tasks include exemplar replay (ER), knowledge distillation (KD), and network expansion (NE). ER-based methods [14, 13] solve the CF problem by storing a small number of old class exemplars. They replay the exemplars with the current session samples to mitigate the old class forgetting when learning new classes. KD-based methods [11, 15, 16] design regularization terms on the image-level feature to preserve previous knowledge when training the model on new data. The NE-based approaches [11, 17] allocate a distinct set of parameters (e.g., tokens, prompts) for each session to learn task-level features. During the training process, parameters associated with old classes irrelevant to the current session are frozen to prevent forgetting old class knowledge.
Despite their promising progress, these existing methods suffer from the feature confusion problem, which hampers their performance in knowledge learning and maintenance, especially when a single image contains both old and new class objects. The feature confusion is two-fold, i.e., the inter-session confusion and intra-feature confusion. Firstly, during the incremental learning process, as the categories of different sessions are disjoint and trained separately, it is difficult to establish sharp decision boundaries between old and new classes, resulting in feature confusion between new and old classes, known as the inter-session confusion [18] problem. This problem is more pronounced in multi-label tasks with multiple targets in a single, leading to a significant performance decline in MLCIL. Secondly, as shown in Fig. 1 (a), all existing methods extract an image-level feature for an image during the incremental learning process. However, as a single image may contain multiple classes in MLCIL, the image-level feature will lead to intra-feature confusion during the incremental learning process. For example, in Fig. 1 (a), the image contains a man riding his bike behind a car. Extracting an image-level feature for this image will not only mix the features of these three classes but also lead to label conflict [11] as the classes are learned in different incremental sessions, resulting in class confusion and semantic conflict within the image-level feature, i.e., the intra-feature confusion problem.
To bridge this gap, we propose a novel framework for MLCIL, named CLass-Independent INcrement (CLIN). Different from the existing methods that extract an image-level feature for MLCIL, we propose a novel class-independent incremental network (CINet) to extract multiple class-level embeddings for input images containing multiple classes and learn and maintain the knowledge of different classes by constructing class-specific tokens as shown in Fig. 1 (b). The CINet comprises a class-level cross-attention module and a class-independent classifier. The former module takes class-specific tokens as input, where each token corresponds to a specific class. It leverages attention mechanisms to extract and generate class-level embeddings associated with each class, effectively producing class-level features and thereby transforming the multi-label problem into multiple single-label problems. Moreover, it can effectively extract local discriminative features and explore correlations between labels as a special Transformer decoder [19, 20, 21]. Subsequently, the class-independent classifier replaces the previous joint classifier, establishing a one-to-one correspondence between class-level embeddings and class labels. It predicts the presence of the corresponding labels through binary classification, reducing interference in learning between new and old classes. During incremental learning, we dynamically extend the class-specific tokens and the class-independent classifier to acquire new knowledge and freeze the tokens and classification heads associated with old classes to prevent forgetting. The proposed CINet not only effectively generates and processes class-level features to address the feature confusion problems but also alleviates inter-session confusion by maintaining knowledge for each old class through a small number of parameters independent of the shared model.
@@
On this basis, to further address the feature confusion problem in MLCIL, we introduce two novel loss functions to optimize the learning of class-specific tokens and class-level embeddings, respectively. Concretely, we address the challenge of applying metric learning to multi-label tasks and propose a contrastive loss on class-level embeddings, i.e. multi-label contrastive loss, aimed at minimizing intra-class distances and maximizing inter-class distances. Moreover, we introduce a token orthogonal loss on the task-specific tokens to alleviate the conflict between old knowledge maintenance and new knowledge learning. Our model achieves significant performance improvements on both existing MLCIL protocols [11, 12]. Moreover, our method maintains high performance in a rehearsal-free setting, significantly saving storage space.
@@
  • We propose a multi-label contrastive loss and introduce a token orthogonal loss to mitigate the interaction between old and new knowledge.
  • We conducted extensive experiments on two well-established protocols, achieving improvements of up to 4.3% and 8.4%, respectively. Furthermore, we evaluated our method in a rehearsal-free setting, achieving state-of-the-art performance and highlighting its robustness.
II Related Work
II-A Single-Label Incremental Learning
@@
Rehearsal-based methods maintain a set of exemplars representing existing data for simultaneous training with new data from the current task. These methods primarily address the challenge of CIL [6, 22, 23, 24] problem. For instance, ER [25] establishes a memory buffer to store samples from previous tasks for retraining with new data, while DER++ [24] suggests applying knowledge distillation penalties to the stored data. Moreover, iCaRL [6] and its variations  [26, 23, 27, 22, 9] prevent forgetting by utilizing the herding technique to select exemplars and formulate distinct KD losses. BIC [28] and other methods [29, 30] incorporate an additional bias correction process to adapt the classification layer. Recent approaches [31, 32, 33] introduce adaptive aggregation networks or emulate the feature space distribution of Oracle to enhance rehearsal-based methods.
Regularization-based methods incorporate a regularization term into the loss function to preserve prior knowledge in updated parameters. 1) Parameter regularization involves minimizing the variance of parameters associated with previous tasks  [34, 35, 36]. For instance, EWC [34] utilizes a Fisher matrix to protect crucial parameters from past tasks. Subsequently, oEWC [36] and other techniques  [35, 37] continuously refine the calculation of parameter significance. 2) Data regularization consolidates previous knowledge by leveraging earlier models as soft teachers [38, 39]. LWF [38], for example, employs knowledge distillation to mitigate forgetting.
@@
Architectural-based methods provide distinct parameters for each task to prevent forgetting. For example, Abati et al. [40, 41] and Rusu et al. [42] have proposed various strategies to isolate old and new task parameters, including duplicating a new network for each task to transfer previous knowledge through lateral connections. Recent architectural approaches [43, 44] combined with the rehearsal methods achieve a better anti-forgetting effect. These methods dynamically expand or prune the network parameters to accommodate the new data at the expense of limited scalability.
@@
Prompt-based methods learn a small set of insertable pre-trained model prompts instead of directly modifying the encoder parameters. For example, L2P [7] creates only a pool of prompts inserted into the model and matches input images to prompts with image-wise prompt queries. DualPrompt [17] proposes multiple pools of prompts, including G-Prompts and E-Prompts for learning task-invariant and task-specific knowledge. Moreover, S-Prompts [45] learns the prompts domain by domain and incrementally inserts the learned prompts into a pool designed for domain-incremental learning. This setting involves learning the same set of classes under covariate distribution task shifts. It is distinct from the CIL setting presented in our paper, where the goal is to learn emerging object classes in new sessions.
II-B Other Incremental Learning
Incremental Object Detection (IOD) focuses on applying incremental learning techniques specifically to object detection. Both knowledge distillation (KD) and experience replay (ER) have found applications in IOD tasks across different detectors. The work by [46] pioneers the use of KD on the output of Faster R-CNN, and subsequent methods [47, 48] extend this by adding KD terms on intermediate feature maps and region proposal networks or by maintaining a set of exemplars for fine-tuning the model. In addition to being implemented on CNN detectors, ER and KD techniques have also been employed in the transformer network DETR [49].
Incremental Semantic Segmentation (ISS) methods can be categorized into regularization-based and replay-based approaches. The former, exemplified by methods like SDR [50] and PLOP [51], introduce various KD strategies to regularize a current model within a latent feature space. The latter approaches [52, 53] rely on an ER strategy involving the retention of a small set of exemplars or pseudo information for previous categories.
Methods developed for incremental object detection (IOD) and incremental semantic segmentation (ISS) cannot be directly applied to MLCIL tasks, as they rely on specific detection and segmentation frameworks or require pixel-level and additional bounding box information. Therefore, research focusing on MLCIL with only image-level annotations holds significant value and importance.
II-C Multi-label Classification
Multi-label classification (MLC) aims to gain a comprehensive understanding of objects and concepts in an image and the proposed methods can be categorized into two main directions:
Label dependencies: Due to the multi-label nature of the data, the co-occurrence relationships between labels can be mined as prior knowledge.  [54] first proposes a ranking-based learning strategy to train CNN for multi-label classification.  [55] and [56] exploit label dependencies by utilizing recurrent neural networks (RNNs) and long short-term memory (LSTM). Moreover, the GCN is also proven to be more effective in modeling label dependencies [57, 58, 2]. Recently, Transformers have also been developed for multi-label classification tasks and have shown great potential [20, 21, 19].  [20] leverages Transformers to exploit the complex dependencies among visual features and labels.  [19] leverages Transformer decoders to adaptively extract local discriminative features for different classes and query the existence of a class object.
Loss functions: The primary challenge in MLC is the significant imbalance between positive and negative instances. To address this, [59] suggests a distribution-balanced loss to reduce imbalance. Another common solution is focal loss [60], which pays more attention to hard samples while less to easy samples. The recent method [61] proposes a novel asymmetric loss, which applies the dynamic down-weight and the hard threshold to easy negative samples while discarding possibly mislabeled samples.
Some MLC methods [20, 21, 19] utilize Transformer decoders to query the existence of class labels, establishing a relationship between queries and labels. However, these methods establish the relationship only once in a supervised learning setting, lacking scalability. These methods face severe catastrophic forgetting in MLCIL, as observed in experiments like FT.
II-D Multi-Label Stream Learning.
@@ II-C Multi-label Classification
Multi-label stream learning (MSL) [62, 63, 64, 65, 66], also known as dynamic multi-label learning or new label emerging, involves reusing previously learned models and creating new classifiers for newly emerging labels. In instance data streams, aside from the initial training set, subsequent data does not have class labels available. Therefore, detecting and modeling new labels becomes the primary challenge. For example, MLF [62] assumes that new labels are linear combinations of other labels and inherits the relationships between labels through classifiers. MuENL [64] and  [67] design detectors based on input features and predicted label attributes and propose algorithms for updating the classifier. DSDL [65] employs streaming label mapping to extract relationships between labels and uses knowledge distillation to transfer knowledge for modeling new labels. FSSL [68] generates label-specific features and effectively fuses these features for newly arrived labels. CIFDM [69] focuses on label correlation and feature embedding space to continuously learn how to predict new labels. ANT [70] trains a student by learning from each teacher’s partial knowledge of label dependencies to infer global dependencies across all labels.
@@
MSL can be considered a special form of weakly supervised learning [64]. Since the data stream for new classes is unlabeled, the core focus is on detecting the emergence of new labels and modeling the relationships between new and known labels rather than on whether known labels are forgotten. This is fundamentally different from the MLCIL setting. MLCIL tasks require learning new classes without forgetting old ones and being able to perform MLC on all previously learned classes. Therefore, MSL approaches cannot be directly applied to MLCIL tasks.
II-E Multi-Label Class Incremental Learning
Multi-label class incremental learning aims to continually construct a unified classifier that can constantly learn and integrate the knowledge and make a comprehensive understanding of class objects in an image. For example, AGCN and its variant [15, 16] employed an augmented Graph Convolutional Network to capture the co-occurrence correlation among classes and incrementally expanded the graph during the learning process. PRS [14] and OCDM [13] enhanced the sampling methods for online data streams, aiming to sample a buffer that is closer to the real distribution to mitigate the imbalance problem. However, these methods belong to online learning settings (i.e., each batch of data can only be accessed once) which exhibit suboptimal performance in offline settings [11].
@@ II-E Multi-Label Class Incremental Learning
Then, KRT [11] defines a more promising task: multi-label class-incremental learning and utilizes an extensible ICA module with token distillation loss to store and transfer knowledge, mitigating catastrophic forgetting. Moreover, APPLE [12] introduces a novel MLCIL benchmark protocol and presents a novel framework that addresses catastrophic forgetting in MLCIL tasks through an adaptive pseudo-label strategy, a cluster sampling method, and a class attention decoder. We build upon the foundation of this method and conduct comprehensive comparative experiments on two offline MLCIL protocols.
III Method

### v2

@@ Abstract
Current research on class-incremental learning primarily focuses on single-label classification tasks. However, real-world applications often involve multi-label scenarios, such as image retrieval and medical imaging. Therefore, this paper focuses on the challenging yet practical multi-label class-incremental learning (MLCIL) problem. In addition to the challenge of catastrophic forgetting, MLCIL encounters issues related to feature confusion, encompassing inter-session and intra-feature confusion. To address these problems, we propose a novel MLCIL approach called class-independent increment (CLIN). Specifically, in contrast to existing methods that extract image-level features, we propose a class-independent incremental network (CINet) to extract multiple class-level embeddings for multi-label samples. It learns and preserves the knowledge of different classes by constructing class-specific tokens. On this basis, we develop a novel loss function, optimizing the learning of class-specific embeddings. The loss aims to distinguish between new and old classes, further alleviating the problem of feature confusion. Extensive experiments on MS-COCO and PASCAL VOC datasets demonstrate the effectiveness of our method for improving recognition performance and mitigating forgetting on various MLCIL tasks.
1 Introduction
@@
In the past decade, deep learning has catalyzed significant breakthroughs in computer vision and multimedia tasks, including image classification [32], semantic segmentation [15], and object detection [4], among others. However, the deep learning-based methods fail to recognize new samples of unseen classes. Directly learning these new samples will lead to catastrophic forgetting (CF), where the model’s performance on historical data significantly deteriorates. In this case, class-incremental learning (CIL) [25, 41] is extensively studied to update the model only with new class data in each task, while preserving the knowledge of old classes. However, in practical applications, an image typically contains multiple objects. For example, photos on social media often contain multiple labels or objects, making traditional CIL methods that assume each image contains only one object unsuitable. Therefore, some work explores studying CIL in the multi-label scenario, named multi-label class-incremental learning (MLCIL) [5, 33, 16]. It aims to gradually learn new classes from label-incomplete multi-label data streams in different learning stages while maintaining performance on old labels.
Compared to the traditional (i.e. single-label) CIL problem, it is more challenging as a single image may contain multiple objects of both the old and new classes, with only new class objects (one or multiple categories) annotated in each learning session. To this end, prior methods to address forgetting for MLCIL tasks include exemplar replay (ER), knowledge distillation (KD), and network expansion (NE). ER-based methods [16] solve the CF problem by storing a small number of old class exemplars. They replay the exemplars with the current session samples to mitigate the old class forgetting when learning new classes. KD-based methods [5, 11] design regularization terms on the image-level feature to preserve previous knowledge when training the model on new data. The NE-based approaches [5, 10, 40] allocate a distinct set of parameters (e.g., tokens, prompts) for each session to learn task-level features. During the training process, parameters associated with old classes irrelevant to the current session are frozen to prevent the forgetting of old class knowledge.
Despite their promising progress, these existing methods suffer from the feature confusion problem, which hampers their performance in knowledge learning and maintenance, especially when a single image contains both old and new class objects. The feature confusion is two-fold, i.e., the inter-session confusion and intra-feature confusion. Firstly, as the categories of different sessions are disjoint and trained separately, it is difficult to establish sharp decision boundaries between old and new classes, resulting in feature confusion between new and old classes, known as the inter-session confusion [24] problem. This problem is more pronounced in multi-label tasks with multiple targets in a single, leading to a significant performance decline in MLCIL. Secondly, as shown in Fig. 1 (a), all existing methods extract an image-level feature for an image during the incremental learning process. However, as a single image may contain multiple classes in MLCIL, the image-level feature will lead to intra-feature confusion during the incremental learning process. For example, in Fig. 1 (a), the image contains a man riding his bike behind a car. Extracting an image-level feature for this image will not only mix the features of these three classes but also lead to label conflict [5] as the classes are learned in different incremental sessions, resulting in class confusion and semantic conflict within the image-level feature, i.e., the intra-feature confusion problem.
To bridge this gap, we propose a novel framework for MLCIL, named CLass-Independent INcrement (CLIN). Different from the existing methods that extract an image-level feature for MLCIL, we propose a novel class-independent incremental network (CINet) to extract multiple class-level embeddings for input images containing multiple classes and learn and maintain the knowledge of different classes by constructing class-specific tokens, as shown in Fig. 1 (b). The CINet comprises a class-level cross-attention module and a class-independent classifier. The former module takes class-specific tokens as input, where each token corresponds to a specific class. It leverages attention mechanisms to extract and generate class-level embeddings associated with each class, effectively producing class-level features, and thereby transforming the multi-label problem into multiple single-label problems. Moreover, it can effectively extract local discriminative features and explore correlations between labels as a special Transformer decoder [21]. Subsequently, the class-independent classifier replaces the previous joint classifier, establishing a one-to-one correspondence between class-level embeddings and class labels. It predicts the presence of the corresponding labels through binary classification, reducing interference in learning between new and old classes. During the incremental learning, we dynamically extend the class-specific tokens and the class-independent classifier to acquire new knowledge and freeze the tokens and classification heads associated with old classes to prevent forgetting. The proposed CINet not only effectively generates and processes class-level features to address the feature confusion problems but also alleviates inter-session confusion by maintaining knowledge for each old class through a small number of parameters independent of the shared model.
@@
On this basis, to further address the feature confusion problem in MLCIL, we introduce two novel loss functions to optimize the learning of class-specific tokens and class-level embeddings, respectively. Concretely, we address the challenge of applying metric learning to multi-label tasks and propose a contrastive loss on class-level embeddings, i.e. multi-label contrastive loss, aimed at minimizing intra-class distances and maximizing inter-class distances. Our model achieves significant performance improvements on both existing MLCIL protocols [5, 33]. Moreover, our method maintains high performance in a more challenging setting [11], significantly saving storage space.
@@
  • We propose a multi-label contrastive loss to further mitigate the interaction between old and new knowledge.
  • Our method exhibits competitive performance comparable to the state-of-the-art methods under two established protocols. Furthermore, in a more challenging setting, it surpassed prior SOTA methods by 8.3% and 4.4% on different benchmarks, respectively, demonstrating superior robustness.
2 Related Work
2.1 Single-Label Incremental Learning
@@
Rehearsal-based methods maintain a set of exemplars representing existing data for simultaneous training with new data from the current task. These methods primarily address the challenge of CIL [25, 35, 6] problem. For instance, ER [27] establishes a memory buffer to store samples from previous tasks for retraining with new data, while DER++ [2] suggests applying knowledge distillation penalties to the stored data. Moreover, iCaRL [25] and its variations  [14, 6, 35] prevent forgetting by utilizing the herding technique to select exemplars and formulating distinct KD losses. BIC [42] and other methods [48] incorporate an additional bias correction process to adapt the classification layer. Recent approaches [1] introduce adaptive aggregation networks or emulate the feature space distribution of Oracle to enhance rehearsal-based methods. CAP [38] stabilizes the feature distribution of previously learned classes by generating augmented prototypes and applying multiple distillation constraints, thereby improving unbiased class-incremental recognition performance.
Rehearsal-Free methods represent a stricter setting, where no samples, prototypes, or class statistics from previous tasks are stored. Early rehearsal-free methods primarily rely on regularization to preserve previously learned knowledge. Parameter regularization minimizes updates to weights associated with earlier tasks [17, 46, 30]. EWC [17] constrains important parameters using the Fisher information matrix, while oEWC [30] and SI [46] further refine the estimation of parameter importance over time to better maintain stability. In contrast, data regularization retains old knowledge by treating previous models as soft teachers [18]. For example, LWF [18] leverages knowledge distillation to mitigate forgetting without storing samples, enabling continual learning in a rehearsal-free setting. More recently, rehearsal-free methods have shifted from purely regularization-based strategies toward more sophisticated mechanisms. BUILD [13] introduces an out-of-distribution (OOD) detection framework that selects pseudo-samples from new-task data to substitute for previously learned classes, enabling fully buffer-free class-incremental learning. DCC-POC [34] improves upon single one-class classifier approaches by employing discriminative and comparable parallel one-class classifiers to enhance both separability and consistency across tasks. SSL-CIL [22] further advances this direction by integrating self-supervised pre-training with prototype learning, effectively mitigating catastrophic forgetting through the decoupling of representation drift, feature confusion, and classifier distortion.
@@
Architectural-based methods provide distinct parameters for each task to prevent forgetting. For example, Abati et al. [23] and Rusu et al. [29] have proposed various strategies to isolate old and new task parameters, including duplicating a new network for each task to transfer previous knowledge through lateral connections. Recent architectural approaches [44, 7, 43] combined with the rehearsal methods achieve a better anti-forgetting effect. These methods dynamically expand or prune the network parameters to accommodate the new data at the expense of limited scalability.
@@
Prompt-based methods learn a small set of insertable pre-trained model prompts instead of directly modifying the encoder parameters. For example, L2P [41] creates only a pool of prompts inserted into the model and matches input images to prompts with image-wise prompt queries. DualPrompt [40] proposes multiple pools of prompts including G-Prompts and E-Prompts for learning task-invariant and task-specific knowledge. Moreover, S-Prompts [39] learns the prompts domain by domain, and incrementally insert the learned prompts into a pool designed for domain-incremental learning. This setting involves learning the same set of classes under covariate distribution task shifts. It is distinct from the CIL setting presented in our paper, where the goal is to learn emerging object classes in new sessions.
2.2 Multi-Label Stream Learning.
@@
Multi-label stream learning (MSL) [45, 50], also known as dynamic multi-label learning or new label emerging, involves reusing previously learned models and creating new classifiers for newly emerging labels. In instance data streams, aside from the initial training set, subsequent data does not have class labels available. Therefore, detecting and modeling new labels becomes the primary challenge. For example, MLF [45] assumes that new labels are linear combinations of other labels and inherits the relationships between labels through classifiers. MuENL [50] designs detectors based on input features and predicted label attributes and proposes algorithms for updating the classifier.
@@
MSL can be considered a special form of weakly supervised learning [50]. Since the data stream for new classes is unlabeled, the core focus is on detecting the emergence of new labels and modeling the relationships between new and known labels, rather than on whether known labels are forgotten. This is fundamentally different from the MLCIL setting. MLCIL tasks require learning labeled new classes without forgetting old ones and being able to perform MLC on all previously learned classes. Therefore, MSL approaches cannot be directly applied to MLCIL tasks.
2.3 Multi-Label Class Incremental Learning
Multi-label class-incremental learning [5, 10, 11, 47] aims to continually build a unified classifier capable of integrating newly acquired knowledge while retaining a comprehensive understanding of multiple co-existing objects within an image. Early attempts, such as AGCN and its variants [8, 9], leveraged augmented graph convolution to model class co-occurrence relationships and incrementally expand the label graph. PRS [16] and OCDM [19] improved online sampling strategies to construct memory buffers closer to real data distributions, thereby alleviating imbalance; however, these online-learning-oriented methods perform suboptimally in offline settings [5].
@@ 2.3 Multi-Label Class Incremental Learning
KRT [5] formally defined the offline multi-label class-incremental learning task and introduced an extensible ICA module with token distillation to preserve and transfer knowledge, effectively mitigating catastrophic forgetting. APPLE [33] further proposed a benchmark protocol and framework combining adaptive pseudo-labeling, cluster-based sampling, and a class-attention decoder to alleviate forgetting in MLCIL. Building upon this line of research, we conduct comprehensive comparisons under two offline MLCIL protocols.
More recently, CSC [10] addressed task-level partial labeling by reducing false positives arising from overconfident predictions, thereby improving retention of previously learned classes. In parallel, RebLL [11] identified the positive–negative imbalance issue in MLCIL and introduced a dual rebalancing framework that jointly tackles data- and loss-level asymmetry, yielding more stable continual multi-label recognition.
3 Methodology